# OpenStereo: A Comprehensive Benchmark for Stereo Matching and Strong Baseline

### 23 Dec 2023 v3

@@ 4.1 Supported Methods
| RAFT-Stereo [22] | FinalPass | - | |
| RAFT-Stereo [22] | FinalPass | 1.18 | |
@@
4.2 Benchmarking Results
Implementation Details All experiments are conducted using the PyTorch [59] version 1.13(or higher) on a cluster of 8 NVIDIA 3090 GPUs (more details in the Supplementary Material).
For benchmarking, it is critical to ensure that the results are reliable and trustworthy. To achieve this, we conducted our experiments using SceneFlow [12] and KITTI15 [16] datasets. As shown in Table 1, in most cases, the reproduced performances of OpenStero are better than the results reported by the original papers. Regarding the KITTI 2015 dataset, submission constraints led us to limit our leaderboard contributions to reproductions of the widely recognized PSMNet [18] and the latest state-of-the-art IGEV [38], as demonstrated in Table 2. We stand by the integrity of our benchmark, asserting its role in providing a reliable and valid evaluation of stereo matching models. OpenStereo is designed to offer the research community in stereo matching a standardized, comprehensive platform for method assessment. This facility enables meaningful and comparative analyses across various models, thereby serving as an invaluable tool for the innovation and scrutiny of emerging algorithms.
5 Revisit Deep Stereo Matching
With the support of OpenStereo, we can conduct a comprehensive reevaluation of various stereo matching methods, including data augmentation, feature extraction, cost construction, disparity prediction and refinement.
Setting For all experiments in this section, each model was trained 90 epochs using a batch size of 64. AdamW is utilized as the optimizer. We implemented a one-cycle [63] learning rate schedule, setting the peak learning rate at 0.0008.
5.1 Data Augmentation
@@
While five data augmentation techniques are prevalent in stereo matching [22, 38], including random crop, color augmentation, eraser transform, flip, and spatial transform, the empirical efficacy of each method specifically for stereo vision tasks remains an underexplored area in the field. Our work extends beyond these common practices, exploring not only these techniques but also experimenting with two additional methods: BatchRandom and Shift, to further understand their impact on stereo matching performance. BatchRandom denotes the technique applied during model training where the size of the random crop varies for each batch. The Shift involves randomly shifting the right image horizontally by a certain range and then cropping both images to a specified size. As shown in Table 3, most data augmentation strategies, except for the combined use of erase transform and random crop, lead to a decline in the model’s performance metrics of SceneFlow. This is because stereo matching involves pixel-level matching, and these data augmentations (color augmentation, flip, and spatial transform) can affect the alignment of pixels. The erase transform benefits stereo matching by enhancing the model’s capacity to cope with occlusions, missing data, and diverse real-world scenarios. It encourages robust feature learning and generalization, ultimately improving model performance. Furthermore, the combination of RandomCrop (320×736) with color augmentation delivers the most striking improvements in KITTI15 dataset performance, achieving the lowest EPE of 1.61 px and the lowest D1_all of 7.81. Such improvements highlight the critical role of color augmentation in boosting the model’s resilience against the variability of color and lighting that naturally occurs in diverse driving environments.
@@
5.2 Feature Extraction
In Table 4, we delve into an ablation study to examine the impact of different backbones and pre-training methods. Initially, we evaluate different backbones without any pre-training. The backbone of FadNet++ [24], despite its low parameter count of 1.77M, results in an EPE of 0.8633, indicating less precision in disparity estimation. Employing MobilenetV2 100 [38] as a backbone improves the EPE to 0.8227, with a slight increase in computational demand (31.83G FLOPs) and parameters (2.78M). The backbone of PSMNet [18], while the most computationally intensive (130.87G FLOPs) with the highest number of parameters (4.46M), achieves a further reduction in EPE to 0.8068, suggesting that the complexity of the network can contribute to the enhancement of stereo matching performance. Furthermore, we delve into the impact of pre-training on the MobilenetV2 100 backbone. Without pre-training, the EPE remains at 0.8227. However, when pre-trained, we observe a substantial improvement, as the EPE drops to 0.7132. This underscores the value of leveraging large-scale image datasets to enhance feature extraction capabilities before fine-tuning specific tasks like stereo matching. In conclusion, our ablation study highlights the critical role of both the architecture’s complexity and pre-training in refining the stereo matching model’s accuracy. It becomes evident that pre-training substantially enhances the model’s capability to estimate disparities with higher precision.
@@ 5.2 Feature Extraction
5.3 Cost Construction
In Table 5, an ablation study on various cost volume strategies for stereo matching is presented. The study begins with simpler 3D cost volume methods like Difference [44] and Correlation [24], yielding higher EPE of 1.02 and 0.81, respectively, at lower computational costs. This suggests that while efficient, these methods may lack the nuanced disparity capture necessary for complex scenes. The Interlaced8 model, introduced by MobileStereoNet [3], achieves the same EPE comparable to the Gwc8 model. However, its computational expense is substantially higher, with a FLOP count of 126.87G which is significantly larger than that of the Gwc8 model. The group-wise correlation models demonstrate a clear trend: as the channel depth increases, the EPE improves, dropping from 0.7132 to 0.6 as channels expand from 8 to 48. This suggests a richer feature capture with more channels, enhancing disparity estimations. Similarly, the concatenation volume shows improved performance with increased channel depth, with a notable decrease in EPE as channels expand from 24 to 64 channels. The combined volume (G8-C16) offers a more optimal balance between computational load and disparity estimation accuracy, which achieves an EPE of 0.62. G16-C24 and G32-C48 do not result in significant improvement in EPE, which stabilizes at 0.60, despite a dramatic increase in computational load, especially for G32-C48 which demands 1460.80 GFLOPs and has 39.37M parameters. These results highlight the delicate balance between accuracy and computational efficiency in designing cost volumes for disparity estimation. While deeper and combined volumes reduce the EPE, the gains might be marginal compared to the significant increase in computational requirements, raising questions about the practicality of these approaches in resource-constrained environments.
@@ 5.3 Cost Construction
| ArgMin | None | 26.37G | 2.69M | 0.7584 |
| ArgMin | RGBRefine | 53.15G | 2.81M | 0.7187 |
| Context | None | 31.83G | 2.78M | 0.7132 |
| Context | RGBRefine | 58.61G | 2.89M | 0.6895 |
| Context | DRNetRefine | 58.57G | 2.89M | 0.6860 |
5.4 Disparity Regression and Refinement
In stereo matching, two prevalent methods for disparity regression are predominantly employed: Differentiable ArgMin [65, 18, 20, 49, 37, 31, 19, 3, 24] and ContextUpsample [62, 22, 38]. The Differentiable ArgMin method, introduced by GCNet [65], calculates disparity by converting matching costs into probabilities via softmax and then computing a weighted sum of these probabilities across all disparity levels. In contrast, the ContextUpsample module, a concept from RAFT [62], focuses on aligning upsampled disparity maps coherently with the image content. This method shows an improvement in EPE to 0.7132, as shown in Table 6, albeit at slightly higher computational costs (31.83G Flops). Furthermore, RGBRefine, a technique involving the concatenation of the left image with the disparity map, further refines the disparity estimation. DRNetRefine [21] slightly improves EPE to 0.6860, maintaining a similar level of computational complexity as RGBRefine.
@@ 5.5.2 Necessity of A Strong baseline
A strong baseline in deep stereo matching research is critical for several key reasons. First, it serves as a vital reference point, enabling a clear assessment of new methods against an established standard. Second, a strong baseline allows for precise evaluation of the impact of specific changes, whether they are new data augmentation methods, different network architectures, or innovative disparity estimation techniques. This helps in isolating and understanding the contribution of each component to the overall performance. Additionally, a solid baseline ensures fair and meaningful comparisons across studies, providing a common ground for evaluating different research outcomes. This is crucial for maintaining consistency and validity in comparative analyses.
@@ 6 A Strong Pipeline: StereoBase
In light of our comprehensive analysis, the goal of this section is to establish a solid yet straightforward baseline model that not only matches but potentially surpasses existing standards in performance. StereoBase, our proposed solution, embodies this objective. It is crafted to be structurally streamlined yet experimentally potent and empirically resilient, setting a new benchmark for future explorations in the field.
@@ 6.1 Pipeline
As shown in Figure 4, we build a simple but strong baseline: StereoBase. StereoBase begins with a carefully curated data preprocessing step. Drawing from our ablation studies, we implement an optimized mix of data augmentation techniques, including RandomCrop, and Erase Transform, proven to enhance model resilience and accuracy. This step ensures that the input data is not only varied and representative of real-world scenarios but also tailored to enhance stereo matching performance. Given the left and the right images, the pre-trained MobileNetV2 [64] networks are used as our foundational backbone, extracting features at a reduced scale of 1/4th the original size to form the cost volume. The cost volume utilizes a combined volume, which combines group-wise correlation volume with an 8-channel and concatenation volume with 16 channels, to ensure an optimal balance between computational load and disparity estimation accuracy. Hourglass networks [38] were implemented for cost aggregation, while context-upample [62] strategies were applied for the final disparity regression. DRNetRefine [21] module was used to refine the final disparity. Training is driven by the smooth L1 [18] loss function. We set the maximum disparity at 192.
@@ 6.2 Comparison with State-of-the-art
In our comprehensive evaluation, we benchmarked StereoBase against current state-of-the-art methods on the Scene Flow test set. This assessment revealed that StereoBase sets a new benchmark in EPE, achieving an unprecedented low score of 0.43. This performance not only establishes a new state-of-the-art but also demonstrates a significant advancement over existing methods, including prominent approaches. The quantitative comparisons, as summarized in Table 7, clearly illustrate the edge of StereoBase in handling complex stereo matching scenarios with greater precision.
@@ 7 Conclusion
This paper introduces OpenStereo, a codebase designed for stereo matching. Our initial endeavor involved the re-implementation of the most state-of-the-art methods within the OpenStereo framework. This comprehensive tool facilitates the extensive reevaluation of various aspects of stereo matching methodologies. Drawing on the insights gained from our exhaustive ablation studies, we proposed StereoBase. StereoBase not only demonstrates the capabilities of our platform but also sets a new standard in the field for future research and development. Through OpenStereo and StereoBase, we aim to contribute a substantial and versatile resource to the stereo matching community, fostering innovation and facilitating more effective and efficient research.

### 08 Mar 2024 v4

@@ 4.1 Recheck on Previous Methods
For benchmarking, it is critical to ensure that the results are reliable and trustworthy. To achieve this, we conducted our experiments using SceneFlow [28] and KITTI2015 [29] datasets. As shown in Table 1, in most cases, the reproduced performances of OpenStero are better than the results reported by the original papers. Regarding the KITTI2015 dataset, submission constraints led us to limit our leaderboard contributions to reproductions of the widely recognized PSMNet [5] and the latest state-of-the-art IGEV [51], as demonstrated in Table 2. We stand by the integrity of our benchmark, asserting its role in providing a reliable and valid evaluation of stereo matching models. OpenStereo is designed to offer the research community in stereo matching a standardized, comprehensive platform for method assessment. This facility enables meaningful and comparative analyses across various models, thereby serving as an valuable tool for the innovation and scrutiny of emerging algorithms.
4.2 Denoising Stereo Matching
With the support of OpenStereo, a comprehensive reevaluation of various stereo matching methods is conducted, including data augmentation, feature extraction, cost construction, disparity prediction and refinement.
@@ 4.2 Denoising Stereo Matching
4.2.1 Data Augmentation
While five data augmentation techniques are prevalent in stereo matching [25, 51], including random crop, color augmentation, eraser transform, flip, and spatial transform, the empirical efficacy of each method specifically for stereo vision tasks remains an underexplored area in the field. Our work extends beyond these common practices, exploring not only these techniques but also experimenting with two additional methods: BatchRandom and Shift, to further understand their impact on stereo matching performance. BatchRandom denotes the technique applied during model training where the size of the random crop varies for each batch. The Shift involves randomly shifting the right image horizontally by a certain range and then cropping both images to a specified size. As shown in Table 3, most data augmentation strategies, except for the combined use of erase transform and random crop, lead to a decline in the model’s performance metrics of SceneFlow. This is because stereo matching involves pixel-level matching, and these data augmentations (color augmentation, flip, and spatial transform) can affect the alignment of pixels. The erase transform benefits stereo matching by enhancing the model’s capacity to cope with occlusions, missing data, and diverse real-world scenarios. It encourages robust feature learning and generalization, ultimately improving model performance. Furthermore, the combination of RandomCrop(320×736) with color augmentation delivers the most striking improvements in Kitti2015 dataset performance, achieving the lowest EPE of 1.61 px and the lowest D1_all of 7.81. Such improvements highlight the critical role of color augmentation in boosting the model’s resilience against the variability of color and lighting that naturally occurs in diverse driving environments.
@@ 4.2.1 Data Augmentation
4.2.2 Feature Extraction
In Table 4, we delve into an ablation study to examine the impact of different backbones and pre-training methods. Initially, we evaluate different backbones without any pre-training. The backbone of FadNet++ [46], despite its low parameter count of 1.77M, results in an EPE of 0.8633, indicating less precision in disparity estimation. Employing MobilenetV2 100 [51] as a backbone improves the EPE to 0.8227, with a slight increase in computational demand (31.83G flops) and parameters (2.78M). The backbone of PSMNet [5], while the most computationally intensive (130.87G flops) with the highest number of parameters (4.46M), achieves a further reduction in EPE to 0.8068, suggesting that the complexity of the network can contribute to the enhancement of stereo matching performance. Furthermore, we delve into the impact of pre-training on the MobilenetV2 100 backbone. Without pre-training, the EPE remains at 0.8227. However, when pre-trained, we observe a substantial improvement, as the EPE drops to 0.7132. This underscores the value of leveraging large-scale image datasets to enhance feature extraction capabilities before fine-tuning specific tasks like stereo matching. In conclusion, our ablation study highlights the critical role of both the architecture’s complexity and pre-training in refining the stereo matching model’s accuracy.
@@ 4.2.2 Feature Extraction
4.2.3 Cost Construction
In Table 5, an ablation study on various cost volume strategies for stereo matching is presented. The study begins with simpler 3D cost volume methods like Difference [21] and Correlation [46], yielding higher EPE of 1.02 and 0.81, respectively, at lower computational costs. This suggests that while efficient, these methods may lack the nuanced disparity capture necessary for complex scenes. The Interlaced8 model, introduced by MobileStereoNet [39], achieves the same EPE comparable to the Gwc8 model. However, its computational expense is substantially higher, with a flop count of 126.87G which is significantly larger than that of the Gwc8 model. The group-wise correlation and concatenation models demonstrate a clear trend: as the channel depth increases, the EPE improves, indicating improved disparity estimations through richer feature capture. The combined volume (G8-C16) offers a more optimal balance between computational load and disparity estimation accuracy, which achieves an EPE of 0.62. G16-C24 and G32-C48 do not result in significant improvement in EPE, which stabilizes at 0.60, despite a dramatic increase in computational load, especially for G32-C48 which demands 1460.80 Gflops and has 39.37M parameters. These results highlight the delicate balance between accuracy and computational efficiency in designing cost volumes for disparity estimation. While deeper and combined volumes reduce the EPE, the gains might be marginal compared to the significant increase in computational requirements, raising questions about the practicality of these approaches in resource-constrained environments.
@@ 4.2.3 Cost Construction
| ArgMin | None | 26.37G | 2.69M | 0.76 |
| ArgMin | RGBRefine [52] | 53.15G | 2.81M | 0.72 |
| ArgMin | Context | 31.83G | 2.78M | 0.71 |
| ArgMin | Context+RGBRefine [52] | 58.61G | 2.89M | 0.69 |
| ArgMin | Context+DRNetRefine [52] | 58.57G | 2.89M | 0.69 |
| ArgMin | ConvGRU [25, 51] | 1363.71G | 12.51M | 0.46 |
4.2.4 Disparity Regression and Refinement
The Differentiable ArgMin [20, 5, 13, 40, 50, 12, 58, 39, 46] introduced by GCNet [20], calculates initial disparity by converting matching costs into probabilities via softmax and then computing a weighted sum of these probabilities across all disparity levels. As shown in Table 6, in this ablation study on disparity refinement for the SceneFlow test datasets, various strategies show differing impacts on model performance. Without refinement, the model has an EPE of 0.76. RGBRefine and Context methods slightly improve EPE to 0.72 and 0.71, respectively, with a modest increase in computational resources. Combining these methods further reduces EPE to 0.69, indicating marginal benefits from their integration. However, employing ConvGRU refinement leads to a substantial improvement, reducing EPE to 0.46, albeit at a significant cost in computational complexity (1363.71 Gflops) and model size (12.51M). This highlights a trade-off between accuracy improvements and increased computational demands.
@@ 4.3.2 Necessity of A Strong Baseline
A strong baseline in deep stereo matching research is critical for several key reasons. First, it serves as a vital reference point, enabling a clear assessment of new methods against an established standard. Second, a strong baseline allows for precise evaluation of the impact of specific changes, whether they are new data augmentation methods, different network architectures, or innovative disparity estimation techniques. This helps in isolating and understanding the contribution of each component to the overall performance. Additionally, a solid baseline ensures fair and meaningful comparisons across studies, providing a common ground for evaluating different research outcomes. This is crucial for maintaining consistency and validity in comparative analyses. In summary, a strong baseline is essential for meaningful advancements in deep stereo matching, ensuring that new developments are substantial, accurately assessed, and broadly applicable.
@@ 5 A Strong Pipeline: StereoBase
In light of our comprehensive analysis, the goal of this section is to establish a strong baseline model that surpasses existing standards in performance. StereoBase, our proposed solution, embodies this objective.
@@ 5.1 Pipeline
We build a strong baseline: StereoBase. StereoBase begins with a carefully curated data preprocessing step. For SceneFlow, KITTI2012, and KITTI2015, we implement an optimized mix of data augmentation techniques, including RandomCrop, ColorAug, and Erase Transform. This step ensures that the input data is not only varied and representative of real-world scenarios but also tailored to enhance stereo matching performance. Given the left and the right images, the pre-trained MobileNetV2 [48] networks are used as our foundational backbone, extracting features at a reduced scale of 1/4th the original size to form the cost volume. The cost volume utilizes a combined volume, which combines group-wise correlation volume with an 8-channel and concatenation volume with 16 channels, to ensure an optimal balance between computational load and disparity estimation accuracy. Hourglass networks [51] were implemented for cost aggregation, while convGRU [51] strategies were applied for the final disparity regression.
| Method | SceneFlow | KITTI 2012 | KITTI 2015 | DrivingStereo | ||||
| EPE | 3-noc | 3-all | D1-bg | D1-fg | D1-all | EPE | D1-all | |
| STTR [24] | 0.43 | - | - | 1.70 | 3.61 | 2.01 | OOM | OOM |
| CFNet [40] | 1.04 | 1.23 | 1.58 | 1.54 | 3.56 | 1.88 | 0.98 | 1.46 |
| AANet [52] | 0.87 | 1.91 | 2.42 | 1.65 | 3.96 | 2.03 | 2.91 | 15.16 |
| COEX [1] | 0.68 | 1.55 | 1.93 | 1.74 | 3.41 | 2.02 | 1.34 | 2.70 |
| FADNet++ [46] | 0.76 | - | - | 1.99 | 3.18 | 2.19 | 1.44 | 5.15 |
| PSMNet [5] | 1.09 | 1.49 | 1.89 | 1.86 | 4.62 | 2.32 | 1.19 | 2.26 |
| GwcNet [13] | 0.76 | 1.32 | 1.70 | 1.74 | 3.93 | 2.11 | 0.99 | 1.36 |
| Mobilestereo-3D [39] | 0.80 | - | - | 1.75 | 3.87 | 2.10 | 1.06 | 1.61 |
| CascadeStereo [12] | 0.72 | - | - | 1.59 | 4.03 | 2.00 | 1.31 | 2.84 |
| IGEV-Stereo [51] | 0.47 | 1.12 | 1.44 | 1.38 | 2.67 | 1.59 | 1.06 | 1.50 |
| StereoBase(Ours) | 0.34 | 1.00 | 1.26 | 1.28 | 2.26 | 1.44 | 1.15 | 2.19 |
@@ 5.2 Comparison with State-of-the-art
In our comprehensive evaluation, we benchmarked StereoBase against current state-of-the-art methods on SceneFlow, KITTI2012, 2015, and DrivingStereo. On the SceneFlow test set, we achieve a new SOTA EPE of 0.34. The quantitative comparisons, as summarized in Table 7, clearly illustrate the edge of StereoBase in handling complex stereo matching scenarios with greater precision. Further, we submitted our results to the KITTI2012 and 2015 leaderboard, where StereoBase outperformed all published methods across all metrics. On KITTI2015, our StereoBase outperforms IGEV [51] by 9.43% on D1-all metric, respectively. In addition, we evaluate the generalization performance of StereoBase. As shown in Table 8, StereoBase exhibited exceptional performance in a zero-shot setting, achieving state-of-the-art results. This evaluation further validates the adaptability and potential of StereoBase in handling diverse and challenging stereo vision tasks.
| Method | KITTI2012 | KITTI2015 | Middlebury | ETH3D | ||
| D1-all(%) | EPE | D1-all(%) | EPE | bad 2.0(%) | bad 1.0(%) | |
| STTR | 49.72 | 6.80 | 40.26 | 6.16 | OOM | 38.89 |
| PSMNet | 30.51 | 4.68 | 32.15 | 5.99 | 33.53 | 18.02 |
| CFNet | 13.64 | 2.27 | 12.09 | 2.89 | 23.91 | 7.67 |
| AANet | 7.23 | 1.27 | 7.72 | 1.41 | 22.45 | 18.77 |
| Mobilestereo-2D | 18.34 | 2.45 | 21.21 | 2.78 | 34.04 | 13.89 |
| Mobilestereo-3D | 18.96 | 2.79 | 19.69 | 3.40 | 29.32 | 13.71 |
| GwcNet | 23.05 | 2.76 | 25.19 | 3.58 | 29.87 | 14.54 |
| COEX | 12.08 | 1.80 | 11.00 | 2.48 | 25.17 | 11.43 |
| FADNet++ | 11.31 | 1.77 | 13.23 | 2.97 | 24.17 | 25.53 |
| CascadeStereo | 11.83 | 1.83 | 12.03 | 2.69 | 27.27 | 11.68 |
| IGEV | 7.80 | 1.40 | 6.35 | 1.31 | 10.51 | 4.44 |
| StereoBase | 4.85 | 0.99 | 5.35 | 1.18 | 9.76 | 3.12 |
@@ 6 Conclusion
This paper introduces OpenStereo, a codebase designed for deep stereo matching. Our initial endeavor involved the re-implementation of the most state-of-the-art methods within the OpenStereo framework. This comprehensive tool facilitates the extensive reevaluation of various aspects of stereo matching methodologies. Drawing on the insights gained from our exhaustive ablation studies, we proposed StereoBase. StereoBase not only demonstrates the capabilities of our platform but also sets a new standard in the field for future research and development. Through OpenStereo and StereoBase, we aim to contribute a substantial and versatile resource to the stereo matching community, fostering innovation and facilitating more effective and efficient research.